# Video-Skill-CoT: Skill-based Chain-of-Thoughts for Domain-Adaptive Video Reasoning

### v1

@@ Abstract
Daeun Leethanks: Equal contribution.  Jaehong Yoon  Jaemin Cho  Mohit Bansal University of North Carolina at Chapel Hill {daeun,jhyoon,jmincho,mbansal}@cs.unc.edu https://video-skill-cot.github.io/
@@ 1 Introduction
Recent work has extended chain-of-thought (CoT) reasoning Wei et al. (2023); Kojima et al. (2022) to multimodal large language models (MLLMs) for video understanding Fei et al. (2024); Feng et al. (2025); Li et al. (2025); Liu et al. (2025); Zhi et al. (2025). However, most prior approaches rely on fixed, general-purpose reasoning traces that are insensitive to domain-specific skills. Fig. 1 (left) shows a t-SNE van der Maaten and Hinton (2008) plot of embeddings of questions from different video datasets, where questions from the same datasets are strongly clustered as they require shared skills/domains. For example, models pretrained on general corpora such as LLaVA-Video-178K Zhang et al. (2024) often lack the nuanced narrative understanding needed in CinePile Rawal et al. (2024). This limits their ability to generalize to unseen domains or specialized skills.
To address this, we propose Video-Skill-CoT (aka Video-SkoT), a novel video understanding framework for creating and leveraging skill-aware CoT supervision, helping effective domain adaptation of MLLMs (Sec. 3). As shown in Fig. 1 (Right), Video-SkoT consists of two main components. First, in skill-based CoT annotation (Sec. 3.2), we introduce a method to automatically construct high-quality, skill-conditioned CoT rationales for video QA tasks. Given a training question, we first extract high-level reasoning skill descriptions (e.g., “Determine object location relative to a person’s orientation” and “Inferring emotional state from expressions and body language”), then cluster them into a shared skill taxonomy (Fig. 1 Right-(a)). Then, each question is annotated with its top-K relevant skills and used to generate a multi-step CoT annotation conditioned on these skills (Fig. 1 Right-(b)). This enables diverse and domain-relevant reasoning traces without requiring manual annotation.
Once we have prepared the skill-based CoT annotations, in skill-specific expert learning (Sec. 3.3 and Fig. 1 Right-(c)), we train skill-specialized expert models with multiple LoRAs Hu et al. (2022). Each expert specializes in a specific set of reasoning capabilities, determined by a predefined group of related questions. During inference, the model routes each input to the expert aligned with the most relevant question group.
@@ 3.1 Problem Setup
Given a video $v$ and a question $q$, our objective is to produce both an answer ${a}$ and a reasoning trace $r$ that offers an interpretable, step-by-step justification. Prior work typically uses a single MLLM $f$ to generate these: $\{r;\,a\}{=}f(q,v)$.
In contrast, Video-SkoT decomposes the reasoning process into two stages: First, given $q$, we select the most relevant expert $e\in\{1,\dots,N^{\mathrm{experts}}\}$ based on the set of pre-defined question groups and predicted required skills. Next, a skill-specific expert MLLM $f^{e}$ then generates a skill-guided reasoning trace $r^{s}$ along with the final answer: $\{r^{s};\,a\}{=}f^{e}(q,v)$. We illustrate Video-SkoT in Fig. 1 (right).
This design enables targeted expert learning and adaptation to diverse reasoning skills in a new video domain. In the following, we describe how we automatically construct the skill-based CoT (Sec. 3.2) and how to train MLLMs with the collected skill-based CoT annotations (Sec. 3.3).
@@ 3.2 Skill-based CoT Annotation
Step 1: Skill Description & Clustering (Fig. 1 Right-(a)).
@@
Step 2: Skill-based CoT Collection (Fig. 1 Right-(b)).
@@ 3.3 Skill-specific Expert Learning
As illustrated in Fig. 1 Right-(c), we perform modularized fine-tuning to learn task-specific knowledge for skill-based CoT training. Specifically, we first project all questions in training set $D^{\text{train}}$ into the text embedding space and perform $k$-means clustering (with $k{=}N^{\text{experts}}{=}5$). Unlike step 2 of Sec. 3.2 where $N^{\text{skills}}$ clusters represent the groups of skill descriptions, these $N^{\text{experts}}$ cluster centroids represent the groups of questions. After assigning each training example to its closest $N^{\text{experts}}$, we conduct parameter-efficient training using the corresponding $N^{\text{experts}}$ expert LoRA Hu et al. (2022) modules, ensuring task-specific adaptation while minimizing interference across skills. During test time, we assign each test question by finding the closest question group by finding the closest question embedding centroids.
@@
Figure 2: Comparison of CoT annotations: (a) regular CoT and (b) our skill-based CoT. Additional examples are provided in Appendix Sec. C.
@@ 4.1 Experiment Setups
To obtain text embeddings (of skill taxonomy in Sec. 3.2 and of questions in Sec. 3.3), we use all-mpnet-base-v2 SentenceTransformers Reimers and Gurevych (2019) implementation. We use LLaVa-Video (7B) Zhang et al. (2024) as a backbone model. Additional training details, including the hyperparameters, specific MLLMs and LLMs used at each stage, are provided in the Appendix  Secs. A.2 and A.1.
@@
We experiment with three different video understanding benchmarks with distinct domains: E.T.-Bench Liu et al. (2024) (temporal understanding), VSI-Bench Yang et al. (2024) (spatial understanding), and CinePile Rawal et al. (2024) (movie narrative understanding). For multiple-choice questions, we report the average accuracy. For temporal captioning tasks in E.T.-Bench, we use the benchmark’s official evaluation script. Baseline MLLMs include mPLUG-Owl Ye et al. (2024), Video-ChatGPT Maaz et al. (2023), Video-LLaMA2 Zhang et al. (2023), LLaVa-OneVision Li et al. (2024a), and LLaVa-Video Zhang et al. (2024). Additional details are provided in the Appendix Sec. A.3.
@@ 4.2 Quantitative Evaluation
Fine-tuned E.T. VSI CinePile
@@
mPLUG-Owl 11.87 - 13.93
Video-ChatGPT 13.02 - 15.08
Video-LLaMA2 8.30 - 44.57
LLaVA-Video 19.35 35.60 55.83
LLaVA-Video 20.32 47.45 56.29
@@
Table 1: Evaluation on domain-specific video reasoning benchmarks. All methods are based on an MLLM (7B).
We compare Video-SkoT to recent MLLM baselines on three video understanding benchmarks (E.T-Bench, VSI-Bench, CinePile) with domains and required skills. Table 1 shows that Video-SkoT consistently outperforms all baselines, achieving improvements of $+4.10$, $+5.70$, and $+1.59$ over the fine-tuned version of LLaVA-Video on E.T.-Bench, VSI-Bench, and CinePile, respectively. These results highlight the effectiveness of our modular, expert-driven framework in enabling domain-adaptive CoT video reasoning by leveraging relevant skills.
@@
We compare the impact of two key components in our framework: (1) skill-based CoT reasoning and (2) skill-specific expert modules. As shown in Tab. 2, our full model, combining both components (Top row), achieves the highest performance. Removing either the skill-specific expert modules (2nd row), the skill-based CoT (3rd row), or both components (last row) consistently leads to performance degradation, highlighting their complementary roles: skill-CoT enables structured reasoning, while expert modules bring modular specialization. This synergy proves essential for improving video understanding.
Skill-CoT (Sec. 3.2) Skill-specific Experts (Sec. 3.3) 3-Task Avg.
@@ 4.3 Qualitative Analysis
Fig. 2 compares the different annotated CoTs from the regular CoT and our skill-based CoT. Given a question about which object is closest to the stove, the regular CoT (left) offers a linear, scene-based narration that lacks structure and includes irrelevant details (“Camera first focuses … it then pans to the right …”), making it often harder to extract key spatial information. In contrast, our skill-based CoT starts by identifying relevant skills (e.g., spatial proximity) and breaking the task into focused sub-questions, like comparing the washer and refrigerator.
@@ 5 Conclusion
Our proposed framework demonstrates strong video reasoning capabilities, generating fine-grained, domain-adaptive rationales based on required skills. However, it may still produce occasional inaccuracies or hallucinations Liu et al. (2023); Wang et al. (2024); Zhou et al. (2024) in its text outputs. Additionally, the overall performance is influenced by the underlying pre-trained backbones, namely, the LLM Achiam et al. (2023) and MLLM Georgiev et al. (2024) used. Nonetheless, we highlight that Video-SkoT can benefit further from future advancements in LLM and MLLM backbones.
@@ A.1 Details of skill description & clustering
Skill Description. To extract skill descriptions given the training dataset, we prompt GPT-4111gpt-4-32k with its questions and answers. (The prompt is provided in Fig. 8) Each extracted skill is written as a concise skill phrase (6–12 words), preserving the core visual or temporal reasoning concept. Here, we intentionally exclude audio-based cues (e.g., sound, speech, or music) in this process. Specific object names (e.g., "TV", "sofa", "John") are replaced with generic terms, and vague terms (e.g., "reasoning", "analysis") are avoided to enhance clarity. We also provide the exact name of the skills in Tab. 3.
@@ A.2 Details of skill-based CoT generation
For skill-based CoT generation, we utilize Gemini-2.0 Flash with video input. As illustrated in Fig. 9, we first prompt Gemini-2.0 to identify the relevant skills and generate corresponding sub-questions and answers. Then, we construct step-by-step reasoning based on this output. Finally, we use GPT-4 to filter and verify the reasoning by assessing its relevance to the ground-truth answers using Fig. 11 as a prompt.
@@ A.4 Prompts
In Figs. 8, 9, 10 and 11, we attach prompts for skill-based CoT annotation. We also attach prompt to generate regular CoT in Fig. 12.
@@ Appendix B Additional Quantitative Results
In Tabs. 4, 6 and 5, we additionally report the per-category performance for each dataset. We also include ablation studies comparing regular CoT vs skill-based CoT, and single-LoRA vs multi-LoRA configurations. Video-SkoT, which combines skill-based CoT with multi-LoRA training, consistently outperforms across all datasets, showing particularly strong gains on reasoning-intensive tasks such as Route Planning in VSI-Bench and temporal understanding tasks in CinePile.
@@ C.1 Skill descriptions over different datasets
In Fig. 7, we visualize the skill descriptions for each dataset after performing skill extraction and clustering (Sec. 3.2). To create the visualization, we first obtain text embeddings using SentenceTransformer and compute $N^{\text{skills}}$ cluster centroids. We then apply t-SNE to reduce the dimensionality of the embeddings for visualization purposes. The results highlight that each domain-specific dataset emphasizes different skill sets, though certain skills are shared across datasets. For instance, the skill “Inferring emotional tone from facial expressions and actions” from CinePile is distinct from “Estimating distance between two objects in the video timeline” from VSI-Bench. However, general skills like “Identifying objects or people” appear across multiple datasets. A more detailed list of the extracted skills is provided in Tab. 3.

### v2

@@ Abstract
Daeun Lee1,thanks: Equal contribution.   Jaehong Yoon1,2,∗   Jaemin Cho1   Mohit Bansal1 1UNC Chapel Hill  2Nanyang Technological University {daeun,jmincho,mbansal}@cs.unc.edujaehong.yoon@ntu.edu.sg https://video-skill-cot.github.io/
@@ 1 Introduction
Recent work has extended chain-of-thought (CoT) reasoning Wei et al. (2023); Kojima et al. (2022) to multimodal large language models (MLLMs) for video understanding Fei et al. (2024); Feng et al. (2025); Li et al. (2025); Liu et al. (2025); Zhi et al. (2025). However, most prior approaches rely on fixed, general-purpose reasoning traces that are insensitive to domain-specific skills. Fig.˜1 (left) shows a t-SNE van der Maaten and Hinton (2008) plot of embeddings of questions from different video datasets, where questions from the same datasets are strongly clustered as they require shared skills/domains. For example, models pretrained on general corpora such as LLaVA-Video-178K Zhang et al. (2024) often lack the nuanced narrative understanding needed in CinePile Rawal et al. (2024). This limits their ability to generalize to unseen domains or specialized skills.
To address this, we propose Video-Skill-CoT (aka Video-SkoT), a novel video understanding framework for creating and leveraging skill-aware CoT supervision, helping effective domain adaptation of MLLMs (Sec.˜3). As shown in Fig.˜1 (Right), Video-SkoT consists of two main components. First, in skill-based CoT annotation (Sec.˜3.2), we introduce a method to automatically construct high-quality, skill-conditioned CoT rationales for video QA tasks. Given a training question, we first extract high-level reasoning skill descriptions (e.g., “Determine object location relative to a person’s orientation” and “Inferring emotional state from expressions and body language”), then cluster them into a shared skill taxonomy (Fig.˜1 Right-(a)). Then, each question is annotated with its top-K relevant skills and used to generate a multi-step CoT annotation conditioned on these skills (Fig.˜1 Right-(b)). This enables diverse and domain-relevant reasoning traces without requiring manual annotation.
Once we have prepared the skill-based CoT annotations, in skill-specific expert learning (Sec.˜3.3 and Fig.˜1 Right-(c)), we train skill-specialized expert models with multiple LoRAs Hu et al. (2022). Each expert specializes in a specific set of reasoning capabilities, determined by a predefined group of related questions. During inference, the model routes each input to the expert aligned with the most relevant question group.
@@ 3.1 Problem Setup
Given a video $v$ and a question $q$, our objective is to produce both an answer a{a} and a reasoning trace $r$ that offers an interpretable, step-by-step justification. Prior work typically uses a single MLLM $f$ to generate these: $\{r;\,a\}{=}f(q,v)$.
In contrast, Video-SkoT decomposes the reasoning process into two stages: First, given $q$, we select the most relevant expert $e\in\{1,\dots,N^{\mathrm{experts}}\}$ based on the set of pre-defined question groups and predicted required skills. Next, a skill-specific expert MLLM $f^{e}$ then generates a skill-guided reasoning trace $r^{s}$ along with the final answer: $\{r^{s};\,a\}{=}f^{e}(q,v)$. We illustrate Video-SkoT in Fig.˜1 (right).
This design enables targeted expert learning and adaptation to diverse reasoning skills in a new video domain. In the following, we describe how we automatically construct the skill-based CoT (Sec.˜3.2) and how to train MLLMs with the collected skill-based CoT annotations (Sec.˜3.3).
@@ 3.2 Skill-based CoT Annotation
Step 1: Skill Description & Clustering (Fig.˜1 Right-(a)).
@@
Step 2: Skill-based CoT Collection (Fig.˜1 Right-(b)).
@@ 3.3 Skill-specific Expert Learning
As illustrated in Fig.˜1 Right-(c), we perform modularized fine-tuning to learn task-specific knowledge for skill-based CoT training. Specifically, we first project all questions in training set $D^{\text{train}}$ into the text embedding space and perform $k$-means clustering (with $k{=}N^{\text{experts}}{=}5$). Unlike step 2 of Sec.˜3.2 where $N^{\text{skills}}$ clusters represent the groups of skill descriptions, these $N^{\text{experts}}$ cluster centroids represent the groups of questions. After assigning each training example to its closest $N^{\text{experts}}$, we conduct parameter-efficient training using the corresponding $N^{\text{experts}}$ expert LoRA Hu et al. (2022) modules, ensuring task-specific adaptation while minimizing interference across skills. During test time, we assign each test question by finding the closest question group by finding the closest question embedding centroids.
@@
Figure 2: Comparison of CoT annotations: (a) regular CoT and (b) our skill-based CoT. Additional examples are provided in Appendix Sec.˜C.
@@ 4.1 Experiment Setups
To obtain text embeddings (of skill taxonomy in Sec.˜3.2 and of questions in Sec.˜3.3), we use all-mpnet-base-v2 SentenceTransformers Reimers and Gurevych (2019) implementation. We use LLaVa-Video (7B) Zhang et al. (2024) as a main backbone model. Additional training details including hyperparameters, the specific MLLMs and LLMs used at each stage, as well as results with the Qwen2.5-VL (7B) backbone are provided in Appendix Secs.˜A.2, A.1 and B.2.
@@
We experiment with three different video understanding benchmarks with distinct domains: E.T.Bench Liu et al. (2024) (temporal understanding), VSI-Bench Yang et al. (2024) (spatial understanding), and CinePile Rawal et al. (2024) (movie narrative understanding). For multiple-choice questions, we report the average accuracy. For temporal captioning tasks in E.T.Bench, we use the benchmark’s official evaluation script. Baseline MLLMs include mPLUG-Owl Ye et al. (2024), Video-ChatGPT Maaz et al. (2023), Video-LLaMA2 Zhang et al. (2023), LLaVa-OneVision Li et al. (2024a), and LLaVa-Video Zhang et al. (2024), GPT4o Achiam et al. (2023a) and Gemini 1.5 Flash, Pro Georgiev et al. (2024). Additional details are provided in the Appendix Sec.˜A.3.
@@ 4.2 Quantitative Evaluation
Fine-tuned E.T.Bench VSI CinePile
@@
Closed-source MLLMs
GPT4o Achiam et al. (2023a) 24.69 34.00 56.06
Gemini 1.5 Pro Georgiev et al. (2024) 26.73 45.40 60.12
Gemini 1.5 Flash Georgiev et al. (2024) 27.92 42.10 58.75
Open-source MLLMs (7B)
mPLUG-Owl Ye et al. (2024) 11.87 - 13.93
Video-ChatGPT Maaz et al. (2023) 13.02 - 15.08
Video-LLaMA2 Zhang et al. (2023) 8.30 - 44.57
LLaVA-Video Zhang et al. (2024) 19.35 35.60 55.83
LLaVA-Video Zhang et al. (2024) 20.32 47.45 56.29
@@
Table 1: Evaluation results on domain-specific video reasoning benchmarks.
We compare Video-SkoT to recent MLLM baselines on three video understanding benchmarks (E.T.Bench, VSI-Bench, CinePile) with domains and required skills. Table˜1 shows that Video-SkoT consistently outperforms all baselines, achieving improvements of +4.10+4.10, +5.70+5.70, and +1.59+1.59 over the fine-tuned version of LLaVA-Video on E.T.Bench, VSI-Bench, and CinePile, respectively. These results highlight the effectiveness of our modular, expert-driven framework in enabling domain-adaptive CoT video reasoning by leveraging relevant skills.
@@
We compare the impact of two key components in our framework: (1) skill-based CoT reasoning and (2) skill-specific expert modules. As shown in Tab.˜2, our full model, combining both components (Top row), achieves the highest performance. Removing either the skill-specific expert modules (2nd row), the skill-based CoT (3rd row), or both components (last row) consistently leads to performance degradation, highlighting their complementary roles: skill-CoT enables structured reasoning, while expert modules bring modular specialization. This synergy proves essential for improving video understanding.
Skill-CoT (Sec.˜3.2) Skill-specific Experts (Sec.˜3.3) 3-Task Avg.
@@
Criterion Regular CoT Skill CoT (Ours) $\Delta$ (Ours - Regular)
Correctness 2.88 $\pm$ 1.61 4.97 $\pm$ 0.16 +2.09
Relevance 3.16 $\pm$ 1.55 4.80 $\pm$ 0.59 +1.64
Coherence 2.80 $\pm$ 1.42 4.73 $\pm$ 0.64 +1.93
Table 3: Human evaluation results comparing Regular CoT and Skill-CoT. Scores are reported as mean $\pm$ standard deviation.
Human Evaluation.
We conduct a human evaluation with five researchers who are familiar with the relevant field, where 15 randomly selected questions were assessed by comparing regular CoT and the proposed Skill-based CoT. Each explanation is rated on a 1–5 Likert scale (5 = best, 1 = worst) across three dimensions: Correctness (factual accuracy), Relevance (task appropriateness), and Coherence (clarity and logical flow). As shown in Tab.˜3, Skill-based CoT consistently outperforms regular CoT across all criteria, with substantial gains in correctness, relevance, and coherence, confirming that our method produces explanations that are more accurate, aligned, and easier to follow. These results provide strong evidence that Skill-based CoT produces explanations that are not only more accurate but also more relevant and human-readable.
@@ 4.3 Qualitative Analysis
Fig.˜2 compares the different annotated CoTs from the regular CoT and our skill-based CoT. Given a question about which object is closest to the stove, the regular CoT (left) offers a linear, scene-based narration that lacks structure and includes irrelevant details (“Camera first focuses … it then pans to the right …”), making it often harder to extract key spatial information. In contrast, our skill-based CoT starts by identifying relevant skills (e.g., spatial proximity) and breaking the task into focused sub-questions, like comparing the washer and refrigerator.
Inference rationale comparison
We compare the inference-time rationales generated by LLaVA-Video trained with (a) regular CoT and (b) the proposed skill-based CoT. During inference, we prompt each model with: “Explain the rationale to answer the question and answer the question.” As shown in Fig.˜3, the model trained with regular CoT produces an incorrect reasoning process, ultimately leading to a wrong answer. In contrast, Video-SkoT successfully generates temporally grounded and precise rationales that more effectively support accurate answer generation.
@@ 5 Conclusion
Our proposed framework demonstrates strong video reasoning capabilities, generating fine-grained, domain-adaptive rationales based on required skills. However, it may still produce occasional inaccuracies or hallucinations Liu et al. (2023); Wang et al. (2024); Zhou et al. (2024) in its text outputs. Additionally, the overall performance is influenced by the underlying pre-trained backbones, namely, the LLM Achiam et al. (2023b) and MLLM Georgiev et al. (2024) used. Nonetheless, we highlight that Video-SkoT can benefit further from future advancements in LLM and MLLM backbones.
@@ A.1 Details of skill description & clustering
Skill Description. To extract skill descriptions given the training dataset, we prompt GPT-4111gpt-4-32k with its questions and answers. (The prompt is provided in Fig.˜8) Each extracted skill is written as a concise skill phrase (6–12 words), preserving the core visual or temporal reasoning concept. Here, we intentionally exclude audio-based cues (e.g., sound, speech, or music) in this process. Specific object names (e.g., "TV", "sofa", "John") are replaced with generic terms, and vague terms (e.g., "reasoning", "analysis") are avoided to enhance clarity. We also provide the exact name of the skills in Tab.˜4.
@@ A.2 Details of skill-based CoT generation
For skill-based CoT generation, we utilize Gemini-2.0 Flash with video input. As illustrated in Fig.˜9, we first prompt Gemini-2.0 to identify the relevant skills and generate corresponding sub-questions and answers. Then, we construct step-by-step reasoning based on this output. Finally, we use GPT-4 to filter and verify the reasoning by assessing its relevance to the ground-truth answers using Fig.˜11 as a prompt.
@@ A.4 Prompts
In Figs.˜8, 9, 10 and 11, we attach prompts for skill-based CoT annotation. We also attach prompt to generate regular CoT in Fig.˜12.
@@ Appendix B Additional Quantitative Results
B.1 Per-category performance
In Tabs.˜5, 7 and 6, we additionally report the per-category performance for each dataset. We also include ablation studies comparing regular CoT vs skill-based CoT, and single-LoRA vs multi-LoRA configurations. Video-SkoT, which combines skill-based CoT with multi-LoRA training, consistently outperforms across all datasets, showing particularly strong gains on reasoning-intensive tasks such as Route Planning in VSI-Bench and temporal understanding tasks in CinePile.
B.2 Qwen2.5-VL backbone
Model Appr. Order Rel. Dist. Route Plan Rel. Dir. Abs. Dist. Obj. Count Obj. Size Room Size Avg
Qwen2.5-VL 10.41 35.29 36.11 37.80 20.43 23.98 54.22 33.06 31.41
Qwen2.5-VL (fine-tune) 18.75 34.45 36.11 38.65 21.88 31.74 62.55 47.74 36.48
Single-LoRA + Regular-CoT 20.83 35.29 36.11 39.92 21.73 34.56 63.33 48.06 37.48
Single-LoRA + Skill-CoT 27.08 35.61 41.66 40.12 21.57 35.33 65.77 48.58 39.47
Ours (Multi-LoRA + Skill-CoT) 31.28 37.89 36.50 43.67 28.12 37.48 67.31 43.79 40.76
Table 8: Qwen2.5-VL (7B) Results on VSI-Bench.
We further evaluate Video-SkoT on VSI-Bench using the Qwen2.5-VL (7B) Bai et al. (2025) backbone. As shown in Tab.˜8, Video-SkoT achieves the highest overall performance (40.76 avg), consistently surpassing both regular-CoT and single-LoRA baselines. These results highlight the robustness and effectiveness of Video-SkoT when applied to a different backbone architecture.
Setting Training Data CRD NPA STA TH TEMP Avg
LLaVA-Video - 56.78 58.53 60.31 60.52 43.02 55.83
LLaVA-Video (fine-tune) CinePile 57.74 58.32 60.44 61.05 43.90 56.29
LLaVA-Video (fine-tune) ET-Bench 53.34 57.02 57.83 60.53 40.55 53.85
Single-LoRA + Regular-CoT ET-Bench 56.23 60.47 58.09 63.68 38.66 55.43
Multi-LoRA + Skill-CoT (Ours) ET-Bench 56.02 61.25 58.88 63.33 41.57 56.21
Table 9: Cross-dataset evaluation on CinePile. Source: ET-Bench $\rightarrow$ Target: CinePile. Backbone: LLaVA-Video.
B.3 Cross-dataset generalization
We evaluate cross-domain generalization from ET-Bench (source) to CinePile (target). As shown in Tab.˜9, Video-SkoT achieves the best average performance (56.21) among ET-Bench–trained variants, performing competitively with the CinePile fine-tuned model (56.29) and surpassing the zero-shot baseline (55.83). This highlights the effectiveness of skill-guided reasoning for transfer across domains.
@@ C.1 Skill descriptions over different datasets
In Fig.˜7, we visualize the skill descriptions for each dataset after performing skill extraction and clustering (Sec.˜3.2). To create the visualization, we first obtain text embeddings using SentenceTransformer and compute $N^{\text{skills}}$ cluster centroids. We then apply t-SNE to reduce the dimensionality of the embeddings for visualization purposes. The results highlight that each domain-specific dataset emphasizes different skill sets, though certain skills are shared across datasets. For instance, the skill “Inferring emotional tone from facial expressions and actions” from CinePile is distinct from “Estimating distance between two objects in the video timeline” from VSI-Bench. However, general skills like “Identifying objects or people” appear across multiple datasets. A more detailed list of the extracted skills is provided in Tab.˜4.